# Optimization of foreground moment deprojection for semi-blind CMB polarization reconstruction

### v1

@@ 1 Introduction
In this work, we introduce the optimized cMILC (ocMILC) pipeline, where the choice of moments to deproject, the values of pivot parameters, and the amount of deprojection of each moment is optimized through a fully data-driven approach. Specifically, the optimal number of moments to deproject before paying significant noise penalty is determined depending on the experimental configuration by applying the strategy of the Generalized Needlet ILC (GNILC) method [32]. This method provides a blind diagnosis of the actual complexity of the Galactic foreground emission across the sky and angular scales, yielding maps of the local number of independent foreground degrees of freedom as byproducts. Such a diagnostic tool is valuable for informing the foreground model parameterization and tuning component separation pipelines according to the observed sky complexity and experimental configuration. The new ocMILC framework is tested and validated on a set of complex foreground scenarios using comprehensive simulations of the Probe of Inflation and Cosmic Origins (PICO) mission concept [6]. We choose this experimental configuration as a proof of concept, given its high sensitivity and broad frequency coverage, allowing us to thoroughly assess the robustness of the proposed approach against various thermal dust and synchrotron models.
The paper is organized as follows: In Section 2, we provide a description of the PICO sky simulations involving four different foreground models. Section 3 outlines our strategy, inspired by GNILC, for conducting a data-driven diagnosis of foreground complexity. Based on this diagnostic, Section 4 introduces the ocMILC pipeline for optimizing foreground moment deprojection and presents the results for different Galactic dust scenarios. Section 5 explores the application of the ocMILC methodology when incorporating various polarized synchrotron and anomalous microwave emission (AME) models. Our conclusions are summarized in Section 6.
@@ 2 Sky simulations
The instrumental noise is simulated as white and isotropic Gaussian realizations, with standard deviations in each pixel corresponding to the polarization sensitivities of the baseline PICO configuration as reported in [6]. The combined noise level for this configuration is equal to $0.87$ $\mu K\cdot\text{arcmin}$. In [6, 30], it has been demonstrated that the targeted sensitivity $r<0.0002$ ($95\%$ CL) can be achieved for most foreground scenarios with the so-called Current Best Estimate (CBE) PICO configuration, featuring a combined noise level of $0.61$ $\mu K\cdot\text{arcmin}$. However, for the purpose of our study relying on the more conservative baseline PICO configuration, with a combined noise level of $0.87$ $\mu K\cdot\text{arcmin}$, we estimate a target sensitivity of about $r<(0.87)^{2}/(0.61)^{2}\times 0.0002=0.0004$ ($95\%$ CL). Here, the noise levels are squared, as the tensor-to-scalar ratio parameter is associated with power spectrum rather than map noise. In Appendix A, we show that, by adopting the CBE PICO configuration, we can indeed achieve $r<0.0002$ ($95\%$ CL) with the ocMILC method.
@@
[…]
@@
In summary, four distinct PICO data sets, hereafter referred to as d1, d4, d7, and d12, are created by combining the respective dust model maps with the PySM s1 synchrotron maps and the random realizations of CMB and instrumental noise. For each data set, 50 simulations are generated, each featuring distinct realizations of CMB and instrumental noise. All maps are generated at the HEALPix resolution $N_{\textrm{side}}=512$, and the resulting PICO frequency maps are smoothed to a common resolution of $\textrm{FWHM}=38^{\prime}$ before undergoing component separation analysis.
3 Diagnosis of foreground complexity with GNILC
The complex properties of the polarized Galactic foregrounds at microwave frequencies are not yet fully understood at the anticipated sensitivity levels of future CMB experiments. It has also been shown that any slight mismodelling of the foregrounds can strongly bias the estimated tensor-to-scalar ratio due to huge amplitude discrepancies between foregrounds and CMB B-modes [28]. Furthermore, the integration of foreground emission along the line-of-sight and within the beam by CMB instruments leads to spatial averaging of each foreground spectral parameter, which results in spectral distortions of the expected foreground SEDs [26, 45] and spectral decorrelation of the foreground emission across frequencies [25, 27]. Additional spectral averaging occurs during data processing steps such as spherical harmonic transforms. All these processes effectively augment the list of expected foreground parameters, adding further complexity to the modelling of the foregrounds. While these effects are small compared to the amplitude of the foreground emission, they can be large compared to the amplitude of the primordial CMB $B$-mode polarization, hence can no longer be neglected for accurate component separation.
To avoid making incorrect assumptions about the number of foregrounds and their spectral properties, it is useful to adopt a data-driven approach which can inform us on the actual complexity of foregrounds in sky observations. The Generalized Needlet ILC (GNILC) method [46, 32] offers an avenue, by providing maps of the effective number of independent foreground components across the sky and different angular scales as byproducts, based on the experiment’s sensitivity and frequency coverage. These GNILC byproduct maps, revealing the effective dimension of the foreground subspace, can in turn inform the Constrained Moment ILC (cMILC) method [23] about the optimal number of nulling constraints or moments to deproject without incurring significant noise penalty, depending on the sky area and the range of angular scales. Besides, parametric component separation methods may also benefit from these diagnostic maps to gauge the expected volume of the foreground parameter space depending on the sky regions.
3.1 Methodology
Instead of making spectral assumptions, GNILC relies on statistical decorrelation to estimate the number of independent foreground degrees of freedom and perform component separation. It capitalizes on the fact that while the noise is largely uncorrelated across frequencies, the foreground emission, $f_{\nu}(\hat{n})$, remains strongly correlated, such that it can be decomposed, at any frequency $\nu$, onto a common basis of a few independent templates $\{t_{i}(\hat{n})\}_{i\in[1,m]}$:
$f_{\nu}(\hat{n})=\sum_{i=1}^{m}\alpha_{\nu,i}(\hat{n})\,t_{i}(\hat{n})\,,$ (3.1)
where the number $m$ of independent templates is typically smaller than the number of frequency channels. The effective number $m$ of these independent templates may vary depending on local signal-to-noise variations across the sky and angular scales. The focus is on mapping this number $m$ rather than the templates themselves to provide insight into the complexity of the foreground sky.
@@ 3.1 Methodology
where (assuming CMB temperature units) $c(\hat{n})$ denotes frequency-independent CMB anisotropies, $f_{\nu}(\hat{n})$ is the Galactic foreground emission, $n_{\nu}(\hat{n})$ the instrumental noise, and $\nu=1,\dotsc,N_{\nu}$, with $N_{\nu}$ the number of available frequency channels. From Equation 3.2, the elements of the $N_{\nu}\times N_{\nu}$ data covariance matrix for frequency pairs $(\nu,\nu^{\prime})$ read as
$C(\hat{n})=F(\hat{n})+C_{N}(\hat{n})\,,$ (3.5)
$\tilde{C}(\hat{n})=\left(\hat{C}_{N}^{-1/2}C\hat{C}_{N}^{-1/2}\right)(\hat{n})\simeq\left(\hat{C}_{N}^{-1/2}F\hat{C}_{N}^{-1/2}\right)(\hat{n})+\mathbb{1}\,,$ (3.6)
where the last equality follows from Equation 3.5 and is exact if the estimated nuisance covariance matrix perfectly captures the statistics of the CMB and noise components, and $\mathbb{1}$ denotes the identity matrix.
where all eigenvalues significantly greater than one correspond to the independent eigenmodes of the Galactic emission which locally dominate over the variance of the noise and CMB components. Therefore, in each point of the sky $\hat{n}$, we are able to estimate the number of independent Galactic foreground modes, $m(\hat{n})$, given the sensitivity of the considered experiment through this model-independent approach where we only need to assume the two-point statistics of the noise and CMB components in the polarization data. The $m$ independent foreground eigenmodes are simply collected in the $N_{\nu}\times m$ matrix $U_{F}$.
We note that, in the above treatment, we considered the CMB as a nuisance whose covariance matrix can be predicted. This may not be exactly the case at large angular scales for CMB $B$-modes since the value of the tensor-to-scalar ratio $r$ is unknown. However, this issue can be addressed by considering only the noise in the nuisance covariance matrix ($C_{N}=N$), in which case our method outputs the number of independent modes from both foregrounds and CMB that dominate over the noise:
$m(\hat{n})=m_{\textrm{fgds}}(\hat{n})+m_{\textrm{CMB}}(\hat{n})\,,$ (3.8)
with $m_{\textrm{CMB}}(\hat{n})=0$ or $1$, as the CMB is fully correlated between frequencies. Given the sensitivity of future CMB experiments, we expect the lensing CMB $B$ mode signal to dominate over instrumental noise across the range of observed angular scales (see e.g. figure 1 of [30]). Therefore,
$m(\hat{n})=m_{\textrm{fgds}}(\hat{n})+1\,.$ (3.9)
The GNILC approach thus allows one to perform a blind diagnosis of the complexity of the Galactic foreground emission in the observed sky and to possibly compare it with that expected from various foreground models.
3.2 Needlet frame implementation
We expect the ’foregrounds-to-noise’ ratio, as traced by the whitened covariance matrix $\tilde{C}(\hat{n})$ (Equations 3.6 and 3.7), to vary not only across the sky but also across angular scales. Most likely, $m_{\textrm{fgds}}$ will be higher at large angular scales where diffuse Galactic emission is brighter than CMB and noise by orders of magnitude, while it will be lower at small angular scales where the noise dominates the covariance budget. This reasoning supports an implementation of our diagnostic tool in wavelet domain. The GNILC pipeline has been developed in a framework of specific spherical wavelets called needlets [47, 48]. In such a framework, the data $d_{\nu}(\hat{n})$ are convolved by a set of needlet bandpass filters in harmonic domain in order to form needlet coefficient maps, $\beta^{j}_{\nu}(\hat{n})$, including sky emission from specific ranges of angular scales:
where $a_{\ell m,\nu}$ are the spherical harmonic coefficients of the frequency maps $d_{\nu}(\hat{n})$, $b^{j}_{\ell}$ is the $j^{\rm th}$ needlet bandpass, and $Y_{\ell m}(\hat{n})$ are spherical harmonics.
In this work, we employ cosine-shaped needlet bandpasses defined as
where $\ell_{\rm peak}=0,50,100,200,300$ for $j=1,\ldots,5$. These needlet bandpasses in harmonic domain are plotted in Figure 1, with each of them covering a specific range of multipoles. By construction, these needlet bands satisfy $\sum_{j}(b^{j}_{\ell})^{2}=1$ for all multipoles, so that the inverse needlet transform
retrieves the initial data. Since the needlet bands have compact support in harmonic domain, the needlet coefficient maps, $\beta^{j}_{\nu}(\hat{n})$, are obtained at a HEALPix $N_{\textrm{side}}$ equal to the smallest power of $2$ larger than $\ell^{j}_{\textrm{max}}/2$, where $\ell^{j}_{\textrm{max}}$ is the maximum multipole of the $j^{\rm th}$ needlet band.
The methodology presented above is applied to the multifrequency needlet coefficient maps, $\beta^{j}_{\nu}(\hat{n})$, instead of the original sky maps, $d{\nu}(\hat{n})$. The data and nuisance covariance matrices, $C^{j}(\hat{n})$ and $C_{N}^{j}(\hat{n})$ of Equation 3.5, are thus computed separately for each range of angular scales selected by the needlet bands. This implementation allows us to produce maps of the number of independent foreground modes, $m^{j}_{\textrm{fgds}}(\hat{n})$, across the sky and in different ranges of angular scales.
Figure 2: Maps of the detected number, $m^{j}_{\textrm{fgds}}$, of independent foreground $E$-mode components (left) and $B$-mode components (right) at needlet scale $j=5$. The results are shown for four distinct foreground scenarios, d1 (first row), d4 (second row), d7 (third row), and d12 (fourth row), and they reveal varying degrees of complexity. These diagnostic maps, which partition the sky into clusters of uniform foreground complexity, were generated using the GNILC methodology outlined in Section 3, applied to full-sky frequency maps of PICO $E$- and $B$-modes.
3.3 Diagnostic maps
Maps of $m^{j}_{\textrm{fgds}}(\hat{n})$ for the fifth needlet scale ($j=5$) are shown in Figure 2 for the four distinct data sets (d1, d4, d7, and d12). They provide the effective number of independent foreground components detected across the sky in both in $E$- and $B$ mode fields at this scale, thereby partitioning the sky into clusters of uniform foreground complexity. These diagnostic maps reveal different degrees of complexity across the sky depending on the underlying foreground scenario, even though all considered foreground models are consistent with Planck observations. In particular, d12 and d1 models exhibit a larger number of independent foreground modes detected across more extensive sky regions compared to d4 and d7 models, with indications of higher complexity for d12. These results hold for both $E$- and $B$-modes. The effective lower complexity of d4 with two MBB components for dust compared to the single-MBB d1 model, is attributed to the absence of spatial variations of each MBB spectral index in d4. The highest level of complexity detected for d12 can be attributed to the multi-layer structure of the dust emission in this model. Therefore, with high polarization sensitivity, as in the case of PICO, our approach offers a reliable tool for discriminating between different foreground models. Moreover, the observed decrease in the effective number of independent foreground modes towards high Galactic latitudes aligns with the spatial extent of Galactic emission in the sky and the prevalence of noise in the data at these latitudes. In $E$-modes, we also observe larger values of $m_{\textrm{fgds}}$ than in $B$-modes over larger regions of the sky. This discrepancy arises from the foreground $E/B$ ratio, which exceeds $2$. However, this trend is partially compensated by the even higher CMB $E/B$ ratio, mitigating the discrepancy between $E$- and $B$-modes concerning the number of independent foreground modes.
In Table 1 we report the excursion set of $m_{\textrm{fgds}}$ observed in the four PICO data sets at the five different needlet scales. For each value, the corresponding sky fraction is also reported in parentheses. As expected, the effective number of independent degrees of freedom of the Galactic emission decreases towards smaller angular scales, where the foreground power decreases while the CMB and noise have relatively larger power. Differences among the four data sets are noticeable. The d1 and d12 data sets exhibit larger values of $m_{\textrm{fgds}}$ at the largest angular scales and over larger portions of the sky at intermediate and small angular scales, both in $E$- and $B$-modes. Compared to other data sets, the d12 data set, which incorporates multilayer dust emission, exhibits higher complexity in broader regions of the sky for all needlet scales. Again, in general, higher values of $m_{\textrm{fgds}}$ are found on larger portions of the sky for $E$-modes.
By diagnosing the number of independent modes of the Galactic foreground emission above the noise level, the blind approach outlined in this section can be leveraged to enhance the optimization of CMB component separation. In the following, we exploit this information to optimize the selection of deprojected moments for the Galactic foregrounds within the framework of the semi-blind cMILC component separation method. Besides, the distribution of $m_{\textrm{fgds}}$ values has the potential to inform parametric component separation methods on the expected dimension of the parameter space for the foreground model to be fitted in specific sky regions and at particular angular scales, minimizing the impact of significant statistical uncertainty in the reconstructed maps of different components.
$E$-modes
Model $m_{\textrm{fgds}}$ ($f_{\textrm{sky}}(\%)$)
$j=1$ $j=2$ $j=3$ $j=4$ $j=5$
d1s1 $6$ $5\ (97),6\ (3)$ $3\ (3),4\ (81),5\ (16)$ $3\ (36.8),4\ (63),5\ (0.2)$ $1\ (0.5),2\ (30.3),3\ (46.2),4\ (23)$
d4s1 $5$ $4\ (1.2),5\ (98.8)$ $3\ (54),4\ (31),5\ (15)$ $3\ (82),4\ (18)$ $1\ (2),2\ (30),3\ (63),4\ (5)$
d7s1 $5$ $5$ $3\ (13),4\ (71),5\ (16)$ $3\ (62),4\ (37.9),5\ (0.1)$ $2\ (31.3),3\ (67),4\ (1.7)$
d12s1 $6$ $5\ (11),6\ (89)$ $4\ (70),5\ (26),6\ (4)$ $3\ (18),4\ (80),5\ (3)$ $1\ (1),2\ (27.8),3\ (26.5),4\ (44.5),5\ (0.2)$
$B$-modes
Model $m_{\textrm{fgds}}$ ($f_{\textrm{sky}}(\%)$)
$j=1$ $j=2$ $j=3$ $j=4$ $j=5$
d1s1 $6$ $4\ (13),5\ (87)$ $3\ (4),4\ (95),5\ (1)$ $2\ (11.5),3\ (28),4\ (60.5)$ $1\ (0.7),2\ (68),3\ (16.3),4\ (15)$
d4s1 $5$ $4\ (16),5\ (84)$ $3\ (59),4\ (40),5\ (1)$ $2\ (12),3\ (72),4\ (16)$ $1\ (3),2\ (67),3\ (25.5),4\ (4.5)$
d7s1 $5$ $4\ (13),5\ (87)$ $3\ (17),4\ (82),5\ (1)$ $2\ (12),3\ (53),4\ (35)$ $2\ (71.5),3\ (27),4\ (1.5)$
d12s1 $6\ (82),7\ (18)$ $5\ (17),6\ (83)$ $4\ (77),5\ (23)$ $2\ (9),3\ (15.3),4\ (75),5\ (0.7)$ $1\ (2),2\ (51),3\ (20.4),4\ (26.6)$
4 The optimized cMILC (ocMILC) method
@@ 4.1 Baseline cMILC approach
The cMILC weights are built to deproject some of these foreground moments by adding to NILC nulling constraints against their homogeneous SEDs, while the remaining unconstrained contamination is left to blind variance minimization:
@@
where $\boldsymbol{w}=\{\text{w}_{\nu}^{j}(\hat{n})\}$ represents the vector of frequency-dependent, pixel-dependent, and scale-dependent cMILC weights, $C=\{C_{\nu\nu^{\prime}}^{j}(\hat{n})\}$ denotes the data covariance matrix with the same dependencies as the weights, $\boldsymbol{\partial^{k}}_{\beta_{\textrm{d}}}\boldsymbol{\partial^{n}}_{T_{\textrm{d}}}\boldsymbol{f}_{\textrm{dust}}(\bar{\beta}_{\textrm{d}},\bar{T}_{\textrm{d}})=\{\partial^{k}_{\beta_{\textrm{d}}}\partial^{n}_{T_{\textrm{d}}}f_{\textrm{dust}}(\nu,\bar{\beta}_{\textrm{d}},\bar{T}_{\textrm{d}})\}$ and $\boldsymbol{\partial^{k}}_{\beta_{\textrm{s}}}\boldsymbol{f}_{\textrm{sync}}(\bar{\beta}_{\textrm{s}})=\{\partial^{k}_{\beta_{\textrm{s}}}f_{\textrm{sync}}(\nu,\bar{\beta}_{\textrm{s}})\}$ are the frequency-dependent but pixel-independent dust and synchrotron moment SEDs, and $\boldsymbol{a}_{\textrm{CMB}}=\{a_{\textrm{CMB},\nu}\}$ stands for the CMB SED, which is independent of frequency in CMB temperature units. The number of nulling constraints on moments in Equation 4.5 is only limited by the number of available frequency channels and the sensitivity of the experiment.
@@
Adding constraints to null some of the foreground moments and, thereby, remove systematic biases generally results in an increase of the variance of the noise and unconstrained foreground components, such as non-included higher-order moments. There is thus a trade-off between deprojection and blind variance minimization, where components that significantly contribute to the overall variance in the data may be better suited for variance minimization rather than deprojection. However, in the baseline cMILC approach, there is no inherent procedure for selecting which moments should be prioritized for deprojection and which ones should be left for variance minimization.
To illustrate some limitations of the baseline cMILC approach, we applied both NILC and cMILC pipelines to different PICO data sets as discussed in Section 2, using the needlet frame implementation presented in Section 3.2. For each data set, we considered $50$ simulations with distinct random realizations of the CMB and instrumental noise, and we implemented three versions of cMILC:
  1. 1.
    cMILC-0, where all zeroth-order moments are deprojected.
  2. 2.
  3. 3.
Foreground and noise residuals are calculated by applying the NILC and cMILC weights derived from the specified data set to the foreground and noise inputs of that simulation, serving illustrative purposes. Throughout this paper, to alleviate residual foreground contamination along the Galactic plane, the angular power spectra for d1, d4, and d7 are computed over $f_{\rm sky}=70\%$ of the sky using the publicly available Planck GAL70 mask applied to the maps. For d12, a more conservative masking strategy is employed, by extending the Planck GAL70 mask through smoothing the average d12 foreground residuals with a $5^{\circ}$ beam and thresholding them to achieve a final sky fraction of $f_{\rm sky}=40\%$. Although the d12 mask is constructed from foreground residuals that may not be accessible in real data analysis, it is employed solely for comparing outcomes from different pipelines within the same clean $40\%$ region of the sky. The angular power spectra are computed using the pymaster Python package [^2] https://namaster.readthedocs.io/en/latest/ of the MASTER estimator [57, 58], which facilitates proper deconvolution of masking and beam smoothing effects [59].
The results for PICO data sets d1 and d7 are reported in Figure 3, showing the angular power spectra of foreground and noise residuals when applying NILC, cMILC-0, cMILC-01, and cMILC-012. Residuals from the most constrained version cMILC-012 are exclusively shown for the d1 data set. The comparison is made against the primordial CMB $B$-mode signal across a range of tensor-to-scalar ratio values, $r\in[0.4\times 10^{-3},4\times 10^{-3}]$, where $r=0.4\times 10^{-3}$ represents the actual upper bound at $95\%$ CL targeted by PICO for its baseline instrumental configuration, as adopted in this work (refer to Section 2 for details on rescaling PICO CBE to baseline sensitivity).
The results from Figure 3 lead to several conclusions. For the PICO d1 data set, the cMILC foreground residuals are lower than those of NILC at intermediate and small angular scales. However, on the largest angular scales, the averaging processes significantly distort the foreground SEDs, and the rigid cMILC deprojection fails to adequately suppress these $-\ell$ distortions of the Galactic emission compared to simple variance minimization. As previously demonstrated in [23], imposing the deprojection of more moments in Equation 4.5 results in a penalty to the residual noise budget. This is particularly noticeable when examining the residuals from the application of cMILC-012 (where all zeroth-, first-, and second-order moments are deprojected) to the PICO d1 data set: the CMB solution, derived from the linear combination of 21 frequency channels with nine fully deprojected moments, suffers from a significant increase in noise contamination. Therefore, a compromise in the optimal number of moments to deproject must be found. Moreover, in this baseline cMILC approach, the moments have been projected out in the natural order of the expansion. However, there is no a priori reason to filter out lower-order moments first in the context of minimum-variance methodologies. It may be more effective to leave some major low-order contributions for blind variance minimization, as they are anyway prevailing in the budget of the observed data covariance matrix of Equation 3.3. When considering PICO maps with different dust models, such as d7 (right panel), the cMILC foreground residuals might even appear brighter compared to NILC in some cases (cMILC-01), indicating that the employed moment deprojection may not entirely capture the actual deviations of the dust model from the MBB SED.

### v2

@@ 1 Introduction
In this work, we introduce the optimized cMILC (ocMILC) pipeline, where the choice of moments to deproject, the values of pivot parameters, and the amount of deprojection of each moment is optimized through a fully data-driven approach. Specifically, the optimal number of moments to deproject before paying significant noise penalty is determined depending on the experimental configuration by applying the strategy of the Generalized Needlet ILC (GNILC) method [32]. This method provides a blind diagnosis of the actual complexity of the Galactic foreground emission across the sky and angular scales, yielding maps of the local number of independent foreground degrees of freedom as byproducts. Such a diagnostic tool is valuable for informing the foreground model parameterization and tuning component separation pipelines according to the observed sky complexity and experimental configuration. The new ocMILC framework is tested and validated on a set of complex foreground scenarios using comprehensive simulations of the Probe of Inflation and Cosmic Origins (PICO) mission concept [6]. We choose this experimental configuration as a proof of concept, given its high sensitivity and broad frequency coverage, allowing us to thoroughly assess the robustness of the proposed approach against various thermal dust and synchrotron models. We note that the validation of the ocMILC pipeline is performed under the assumption of perfectly reconstructed frequency maps from time-ordered data. However, as shown by the latest analyses of the Planck data (NPIPE [33], SRoll [34, 35], Cosmoglobe [36]), a parametric model of the Galactic emission may still be needed for the low-level data processing steps, as calibration, bandpass corrections, and multi-component destriping, involved in producing high-fidelity frequency maps.
The paper is organized as follows: In Section 2, we provide a description of the PICO sky simulations, considering four foreground scenarios with different dust models. Section 3 outlines our component separation methodology, first reviewing the basics of the cMILC approach (Section 3.1), then describing the proposed ocMILC pipeline (Section 3.2), which involves a data-driven diagnosis of foreground complexity inspired by GNILC (Section 3.2.1) and subsequent optimization of foreground moment deprojection (Section 3.2.2). Section 4 presents our results for the different foreground scenarios. Our conclusions are summarized in Section 5.
@@ 2 Sky simulations
The instrumental noise is simulated as white and isotropic Gaussian realizations, with standard deviations in each pixel corresponding to the polarization sensitivities of the baseline PICO configuration as reported in [6]. The combined noise level for this configuration is equal to $0.87$ $\mu K\cdot\text{arcmin}$. In [6, 30], it has been demonstrated that the targeted sensitivity $r<0.0002$ ($95\%$ CL) can be achieved for most foreground scenarios with the so-called Current Best Estimate (CBE) PICO configuration, featuring a combined noise level of $0.61$ $\mu K\cdot\text{arcmin}$. However, for the purpose of our study relying on the more conservative baseline PICO configuration, with a combined noise level of $0.87$ $\mu K\cdot\text{arcmin}$, we estimate a target sensitivity of about $r<(0.87)^{2}/(0.61)^{2}\times 0.0002=0.0004$ ($95\%$ CL). Here, the noise levels are squared, as the tensor-to-scalar ratio parameter is associated with power spectrum rather than map noise. In Section 4.2, we show that, by adopting the CBE PICO configuration, we can indeed achieve $r<0.0002$ ($95\%$ CL) with the ocMILC method.
@@
    where $X=\{\textit{Q},\textit{U}\}$ are Stokes parameters, $B_{\nu}(T)$ is the blackbody spectrum, $\beta_{\textrm{d}}$ is the dust spectral index, $T_{\textrm{d}}$ is the dust temperature, and $X_{d}(\hat{n},\nu_{\textrm{d}})$ represents the dust template at a reference frequency $\nu_{\textrm{d}}$. Both $\beta_{\textrm{d}}$ and $T_{\textrm{d}}$ vary across the sky, as estimated from the Planck data analysis using the Commander algorithm [42]. The thermal dust template is given by the Commander dust $Q$ and $U$ maps at $353$ GHz of the Planck $2015$ data release, smoothed with a Gaussian beam of full width at half maximum (FWHM) $2.6^{\circ}$, with stochastic small-scale fluctuations added via the procedure described in [40].
@@
In summary, four distinct PICO data sets, hereafter referred to as d1, d4, d7, and d12, are created by combining the respective dust model maps with the PySM s1 synchrotron maps and random realizations of CMB. At each frequency, these components are first smoothed using a Gaussian beam with the corresponding FWHM of that PICO channel and then co-added to instrumental noise. Such resulting PICO frequency maps are ultimately brought to a common resolution of $\textrm{FWHM}=38^{\prime}$ and downgraded to a HEALPix resolution of $N_{\textrm{side}}=512$ before undergoing component separation analysis. For each data set, 50 simulations are generated, each featuring distinct realizations of CMB and instrumental noise.
3 Component separation methodology
@@
where (assuming CMB temperature units) $c(\hat{n})$ denotes frequency-independent CMB anisotropies, $f_{\nu}(\hat{n})$ is the Galactic foreground emission, $n_{\nu}(\hat{n})$ the instrumental noise, $\hat{n}$ the sky position, and $\nu=1,\dotsc,N_{\nu}$, with $N_{\nu}$ the number of available frequency channels.
An estimate $\hat{c}(\hat{n})$ of CMB anisotropies can be formed via a weighted linear combination of the data:
$\hat{c}(\hat{n})=\boldsymbol{w}^{T}\cdot\boldsymbol{d}=\sum_{\nu}w_{\nu}(\hat{n})\,d_{\nu}(\hat{n})\,,$ (3.2)
where the specific weights $w_{\nu}$ assigned to data $d_{\nu}$ across frequency channels are sought to minimize the variance due to foreground and noise contamination while preserving the CMB signal $c(\hat{n})$. This is the strategy of ILC methods such as NILC [21]. In this section, we seek to optimize these weights to minimize foreground variance rather than global variance as a function of sky regions and angular scales with our proposed ocMILC (optimized cMILC) pipeline. We first review the basics of the cMILC formalism (Section 3.1) which forms the skeleton of the proposed ocMILC pipeline, the specific properties of which are described afterward (Section 3.2).
@@ 3.1 Baseline cMILC approach
The cMILC weights $\boldsymbol{w}$ are built to deproject some of these foreground moments by adding to NILC nulling constraints against their homogeneous SEDs, while the remaining unconstrained contamination is left to blind variance minimization:
@@
where $\boldsymbol{w}=\{w_{\nu}(\hat{n})\}$ represents the vector of frequency-dependent, pixel-dependent cMILC weights, $C=\{C_{\nu\nu^{\prime}}(\hat{n})=\langle d_{\nu}(\hat{n})\,d_{\nu^{\prime}}(\hat{n})\rangle\}$ denotes the data covariance matrix with the same dependencies as the weights, $\boldsymbol{\partial^{k}}_{\beta_{\textrm{d}}}\boldsymbol{\partial^{n}}_{T_{\textrm{d}}}\boldsymbol{f}_{\textrm{dust}}(\bar{\beta}_{\textrm{d}},\bar{T}_{\textrm{d}})=\{\partial^{k}_{\beta_{\textrm{d}}}\partial^{n}_{T_{\textrm{d}}}f_{\textrm{dust}}(\nu,\bar{\beta}_{\textrm{d}},\bar{T}_{\textrm{d}})\}$ and $\boldsymbol{\partial^{k}}_{\beta_{\textrm{s}}}\boldsymbol{f}_{\textrm{sync}}(\bar{\beta}_{\textrm{s}})=\{\partial^{k}_{\beta_{\textrm{s}}}f_{\textrm{sync}}(\nu,\bar{\beta}_{\textrm{s}})\}$ are the frequency-dependent but pixel-independent dust and synchrotron moment SEDs, and ${\boldsymbol{a}_{\textrm{CMB}}=\{a_{\textrm{CMB},\nu}\}}$ stands for the CMB SED, which is independent of frequency in CMB temperature units. The number of nulling constraints on moments in Equation 3.7 is only limited by the number of available frequency channels and the sensitivity of the experiment.
@@
Adding constraints to null some of the foreground moments in Equation 3.7 and, thereby, remove systematic biases generally results in an increase of the variance of the noise and unconstrained foreground components, such as non-included higher-order moments. There is thus a trade-off between deprojection and blind variance minimization, where components that significantly contribute to the overall variance in the data may be better suited for variance minimization rather than deprojection. However, in the baseline cMILC approach, there is no inherent procedure for selecting which moments should be prioritized for deprojection and which should be left for variance minimization, nor is there a method for assessing the optimal number of moments to deproject before incurring significant noise penalties. Moreover, if the effective SED of foreground emission turns out to highly deviate from the adopted baseline SEDs of Equations 3.3 and 3.5, moment expansion may not fully capture such distortions, in which case partial deprojection or variance minimization could be more effective. Therefore, there is room for optimization of moment deprojection in cMILC, as anticipated and preliminarily explored in [23].
3.2 Optimized cMILC (ocMILC) method